Ask
your own question, for FREE!
Mathematics
10 Online
The probability of a person over the age of 50 developing cancer is .06 and the probability that she/he develops heart problems is .09. If a patient over 50 is selected at random, what is the probability that the patient will develop at least one of the diseases? Round your answer to two decimal places.
Still Need Help?
Join the QuestionCove community and study together with friends!
help
P(at least one)=1-P(gets none)=1-P(no C and no H) =1-P(no C)*P(no H) =1-(1-.06)(1-.09)=.1446
This can't be solved without knowing something about the probability that they will get both. The two probabilities are not independent.
I believe that is an assumption to be made (that they are independent) based on previous problems asked by hscheide
Ah, understood.
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
13 hours ago
1 Reply
0 Medals
5 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
3 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal